The graduate selection committee wants to select the top 10% of applicants. On a standardized test with a mean of 500 and a standard deviation of 100, what would be the cutoff score for selecting the top 10% of applicants, assuming that the standardized test is normally distributed?

Answer by stanbon(48568)

The graduate selection committee wants to select the top 10% of applicants. On a standardized test with a mean of 500 and a standard deviation of 100, what would be the cutoff score for selecting the top 10% of applicants, assuming that the standardized test is normally distributed?
---------------------------------------------------
Find the z-value that has right-tail of 10%.
z = invNorm(0.90) = 1.2816
--------------------------------
Find the x-value that corresponds to that z-score:
x = zs + u
x = 1.2816*100 + 500
x = 128.16 + 500
x = 628.16
=================
Cheers,
Stan H.